### v2

Spanning Trees Minimizing Branching Costs ***An extended abstract of some of the results of this paper appeared in [23].
[…]
Department of Computer Science
University of Salerno
Italy
{lgargano, arescigno}@unisa.it
The Minimum Branch Vertices Spanning Tree problem aims to find a spanning tree $T$ in a given graph $G$ with the fewest branch vertices, defined as vertices with a degree three or more in $T$. This problem, known to be NP-hard, has attracted significant attention due to its importance in network design and optimization. Extensive research has been conducted on the algorithmic and combinatorial aspects of this problem, with recent studies delving into its fixed-parameter tractability.
In this paper, we focus primarily on the parameter modular-width. We demonstrate that finding a spanning tree with the minimum number of branch vertices is Fixed-Parameter Tractable (FPT) when considered with respect to modular-width. Additionally, in cases where each vertex in the input graph has an associated cost for serving as a branch vertex, we prove that the problem of finding a spanning tree with the minimum branch cost (i.e., minimizing the sum of the costs of branch vertices) is FPT with respect to neighborhood diversity.
Keywords: Spanning tree, Fixed parameterized algorithms, Modular-width, Neighborhood diversity
Declarations
  • Funding: This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU.
  • Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.
1 Introduction
Let $G=(V,E)$ be an undirected connected graph where $V$ is the set of vertices and $E$ is the set of edges. Given a spanning tree $T$ of $G$, a branch vertex is a vertex having degree three or more in $T$. We denote by $b(G)$ the smallest number of branch vertices in any spanning tree of $G$. We study the following problem:
Minimum Branch Vertices (MBV)
Instance: A connected graph $G=(V,E)$.
Goal: Find a spanning tree of $G$ having $b(G)$ branch vertices.
By noticing that the only tree without branch vertices is a path, we know that $b(G)=0$ if and only if $G$ is Hamiltonian.
The problem of determining a spanning tree with a limited number of branch vertices, though fundamentally a theoretical question, originated from addressing challenges in wavelength-division multiplexing (WDM) technology within optical networks. WDM is a highly effective method for utilizing the full bandwidth potential of optical fibers, thereby meeting the high bandwidth demands of the Internet [28].
Multicasting involves the simultaneous transmission of information from a single source to multiple destinations. In Wavelength-Division Multiplexing (WDM) systems, light trees are crucial for efficiently managing multicast communications by duplicating and routing specific wavelength channels. Light-splitting switches in these systems separate individual wavelengths within an optical signal. Due to the high cost of light-splitting switches, it is essential to minimize the number of nodes equipped with these devices. This need gives rise to the Minimum Branch Vertices (MBV) problem, which seeks to reduce the number of light-splitting switches in a light tree[20]. In Cognitive Radio Networks, as well as in 5G technologies that operate across a broad spectrum of frequencies, managing the switching costs associated with transitions between different service providers is of utmost importance. This is essential not only for minimizing delays but also for optimizing energy consumption [27, 37].
The Minimum Branch Vertices (MBV) problem has been extensively studied from both algorithmic and graph-theoretic perspectives. Most prior research has focused on establishing upper bounds for the number of branch vertices in the resulting tree, though these bounds were often not tight. Gargano et al. [21] proved that determining whether a graph $G$ has a spanning tree with at most $k$ branch vertices is NP-complete, for each $k\geq 1$, even in cubic graphs. In the same paper, they provided an algorithm that finds a spanning tree with one branch vertex if every set of three independent vertices in $G$ has a degree sum of at least $|V(G)|-1$. Salamon [36] developed an algorithm that finds a spanning tree with $O(\log|V(G)|)$ branch vertices for graphs where each vertex has a degree of $\Omega(n)$; moreover, an approximation factor better than $O(\log|V(G)|)$ would imply that $NP\subseteq DTIME(n^{O(\log\log n)})$. Sufficient conditions for a connected claw-free graph to have a spanning tree with $k$ branch vertices are presented in [34].
Integer linear formulations of the MBV problem and its variants are discussed in [9, 10, 11], including different relaxations and numerical comparisons of these relaxations. Hybrid integer linear programs for MBV are considered and solved with branch-and-cut algorithms in [39]. Decomposition methods for solving the MBV problem are explored in [31, 33, 35], while other heuristics are presented in [32, 38, 40].
A complementary formulation called the Maximum Path-Node Spanning Tree (MPN), which aims to maximize the number of vertices with a degree of at most two, is investigated in [12]. The authors prove that MPN is APX-hard and provide an approximation algorithm with a ratio of 6/11. Related gathering processes are considered in [5, 6, 7, 24, 25].
1.1 Graph Partitioning.
In order to address the Minimum Branch Vertices (MBV) problem, we also consider two related graph partitionings that can be of their own interest.
A spider is defined as a tree having at most one branch vertex, which is termed the center of the spider if it exists, otherwise any vertex can serve as the center. A path-spider cover of a graph $G$ consists of one spider and additional vertex-disjoint paths whose union includes every vertex of $G$. We denote by $spi(G)$ the least integer $p$ such that $G$ has a path-spider cover with $p-1$ paths. We define and study the following problem:
Path-Spider Cover (PSC)
Instance: A graph $G=(V,E)$.
Goal: Find a path-spider cover of $G$ with $spi(G)-1$ paths.
We will denote by by ${\cal{P}}_{spi(G)}$ a path-spider cover of $G$ with $spi(G)-1$ paths.
Moreover, we will make use of the Partitioning into Paths problem defined below. A partition of a graph $G$ into paths is a set of (vertex-)disjoint paths of $G$ whose union contains every vertex of $G$. We denote by $ham(G)$ the least integer $p$ such that $G$ has a partition into $p$ paths. Clearly, $spi(G)\leq ham(G)$.
Partitioning into Paths (PP)
Instance: A graph $G=(V,E)$.
Goal: Find a partition of $G$ into $ham(G)$ paths.
We will denote by ${\cal{P}}_{ham(G)}$ a partition of $G$ into $ham(G)$ paths.
Notice that the Partitioning into Paths problem was already considered in [18]; however as originally defined, it only asks for the value $ham(G)$, while we ask for the actual path partitioning of $G$. Recently, the PP problem and some its variants have been considered in [16].
When referring to a path $P$ in $G$, we denote its end-points as $f(P)$ and $s(P)$, distinguishing them as the first and the second end-point of $P$, respectively. Additionally, if $P$ represents a spider in $G$, either $f(P)$ or $s(P)$ is used interchangeably to denote the center of $P$.
1.2 Our Results
In Section 2, we present an FPT algorithm for MBV parameterized by modular-width. To this aim, we also design FPT algorithms for PSC and PP parameterized by modular-width; they are presented in Section 3.
Additionally, in Section 4 we direct our focus to a scenario where the cost associated with selecting a vertex as a branch vertex varies across the graph. Consequently, we delve into the case where each vertex $v\in V$ is assigned a specific non-negative integer weight $w(v)$ and the cost of a spanning tree $T$ is $w(T)=\sum w(u)$, where the sum is over all the branch vertices of $T$. The goal of the weighted version of the MBV is to find a spanning tree $T$ of the input graph that minimizes $w(T)$. For the weighted MBV problem we present an FPT algorithm parameterized by neighborhood diversity. An FPT algorithm for the case of uniform cost was given in [22].
1.3 Parameterized Complexity.
Parameterized complexity is a refinement to classical complexity theory in which one takes into account not only the input size, but also other aspects of the problem given by a parameter $p$. A problem with input size $n$ and parameter $p$ is called fixed parameter tractable (FPT) if it can be solved in time $f(p)\cdot n^{c}$, where $f$ is a computable function only depending on $p$ and $c$ is a constant.
It was recently proven that MBV is FPT when parameterized by treewidth [4]. The algorithm presented in [4] runs in time $O(4^{2{\tt tw}}{\tt tw}^{4{\tt tw}+1}({\tt tw}+1)\;n)$, where tw is the treewidth of the input graph. On the other hand, it was shown in [17] that the problem is $W[1]$-hard when parameterized by clique-width. Specifically, in [17] it was proven that the (MBV special case) hamiltonian path problem is $W[1]$-hard when parameterized by clique-width.
In this paper, we are interested in assessing the complexity of MBV when parameterized by modular width and, in the weighted case, by neighborhood diversity.
These graph parameters could cover dense graphs but still allow FPT algorithms for the problems lost to clique-width. See Figure 1 for a relation among the above parameters.
Figure 1: A summary of the relations holding among some popular parameters. We use mw$(G)$, tw$(G)$, cw$(G)$, nd($G$), and vc$(G)$ to denote modular-width, treewidth, cliquewidth, neighborhood diversity, and minimum vertex cover of a graph $G$, respectively. Solid arrows denote generalization, e.g., modular-width generalizes neighborhood diversity. Dashed arrows denote that the generalization may exponentially increase the parameter.
Modular-width. Modular-width was introduced in [18] and is defined as follows.
Consider graphs that can be obtained from an algebraic expression that uses the following operations:
  • (O1)
    Create an isolated vertex;
  • (O2)
    the disjoint union of 2 graphs denoted by $G_{1}\oplus G_{2}$, i.e., $G_{1}\oplus G_{2}$ is the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})$;
  • (O3)
    the complete join of 2 graphs denoted by $G_{1}\otimes G_{2}$, i.e., $G_{1}\otimes G_{2}$ is the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})\cup\{\{v,w\}:v\in V(G_{1})\;\mbox{and}\;w\in V(G_{2})\}$;
  • (O4)
    the substitution operation with respect to some graph ${{H}}$ with vertex set $\{1,2,\ldots,n\}$ i.e., for graphs $G_{1},\ldots,G_{n}$ the substitution of the vertices of ${{H}}$ by the graphs $G_{1},\ldots,G_{n}$, denoted by ${{H}}(G_{1},\ldots,G_{n})$, is the graph with vertex set $\bigcup_{i=1}^{n}V(G_{i})$ and edge set $\bigcup_{i=1}^{n}E(G_{i})$$\cup\{\{u,v\}\ |\ u\in V(G_{i}),v\in V(G_{j}),\{i,j\}\in E({{H}})\}$. Hence, ${{H}}(G_{1},\ldots,G_{n})$ is obtained from ${{H}}$ by substituting every vertex $i\in V({{H}})$ with the graph $G_{i}$ and adding all edges between the vertices of a graph $G_{i}$ and the vertices of a graph $G_{j}$ whenever $\{i,j\}\in E({{H}})$.
The width of an algebraic expression, that uses only the operations (O1)–(O4), is the maximum number of operands used by any occurrence of the operation (O4). The modular-width of a graph ${{G}}$, denoted ${\tt{mw}}({{G}})$, is the least integer $m$ such that ${{G}}$ can be obtained from such an algebraic expression of width at most $m$.
We recall that an algebraic expression of width ${\tt{mw}}({{H}})$ can be constructed in linear time [41].
Given a graph $G=(V,E)$, denote by $N(u)$ the set of neighbors of vertex $u\in V$. A module of a graph $G$ is a set $M\subseteq V$ such that for all $u,v\in M,\ N(u)\setminus M=N(v)\setminus M.$ Operations (O1)-(O4) taken to construct a graph, form a parse-tree of the graph. A parse tree of a graph $G$ is a tree $T(G)$ that captures the decomposition of $G$ into modules. The leaves of $T(G)$ represent the vertices of $G$ (operation (O1)). The internal vertices of $T(G)$ capture operations on modules: Disjoint union of its children (operation (O2)), complete join (operation (O3)) and substitution (operation (O4)). Figure 2 depicts a graph $G$ and the corresponding parse tree.
Figure 2: (a) A graph $G$. (b) The parse tree $T(G)$ associated with a decomposition of $G$ into modules. The width of the presented decomposition is $5$. (c) A representation of the decomposition of $G$ into modules. (d) An expression describing $G$ in terms of the operations (O1)-(O4).
Given a graph ${{G}}={{H}}(G_{1},\ldots,G_{n})$, we will refer to the graphs $G_{1},\ldots,G_{n}$ also as the modules of ${{G}}$. Notice that given the graph ${{G}}={{H}}(G_{1},\ldots,G_{n})$, by the operations O(1)-(O4), one has that all the vertices of $G_{i}$ share the same neighborhood outside $G_{i}$; indeed,
$\begin{array}[]{ll}{\{\{u,v\}\,|\,u\in V(G_{i}),v\in V(G_{j})\}\subseteq E({{G}})}&\mbox{ if $\{i,j\}\in E({{H}})$}\\ {\{\{u,v\}\,|\,u\in V(G_{i}),v\in V(G_{j})\}\cap E({{G}})=\emptyset}&\mbox{ if$\{i,j\}\notin E({{H}})$}\end{array}$ (1)
for each $i,j=1,\ldots,n$ with $i\neq j$.
Neighborhood diversity. The neighborhood diversity parameter, was first introduced by Lampis in [30]. It has then received much attention [1, 2, 8, 13, 14, 15, 19, 24, 26, 42], also due to the fact that, contrary to other popular parameters, it is computable in linear time.
Given a graph $G=(V,E)$, two vertices $u,v\in V$ are said to have the same type if $N(v)\setminus\{u\}=N(u)\setminus\{v\}$. The neighborhood diversity nd$(G)$ of a graph $G$ is the minimum number ${\tt{nd}}$ of sets in a partition $V_{1},V_{2},\ldots,V_{\tt{nd}}$, of the vertex set $V$, such that all the vertices in $V_{i}$ have the same type, for $i=1,\ldots,{\tt{nd}}$.
The family $\{V_{1},V_{2},\ldots,V_{\tt{nd}}\}$ is called the type partition of $G$. Notice that each $V_{i}$ induces either a clique or an independent set in $G$. Moreover, for each $V_{i}$ and $V_{j}$ in the type partition, either each vertex in $V_{i}$ is a neighbor of each vertex in $V_{j}$ or no vertex in $V_{i}$ has a neighbor in $V_{j}$. Hence, between each pair $V_{i}$ and $V_{j}$ there is either a complete bipartite graph or no edge at all. Starting from a graph $G$ and its type partition ${\cal{V}}=\{V_{1},\ldots,V_{\tt{nd}}\}$, we can see each element of ${\cal{V}}$ as a vertex of a new graph $H$, called the type graph of $G$, with
- $V(H)=\{1,2,\cdots,{\tt{nd}}\}$
- $E(H)=\{\{x,y\}\ |\ x\neq y\mbox{ and for each $u\in V_{x}$, $v\in V_{y}$ it holds that $\{u,v\}\in E(G)$ }\}$.
Hence, a graph $G$ of neighborhood diversity ${\tt{nd}}$ can be seen as $G=H(G_{1},\ldots,G_{\tt{nd}})$, in which $G_{i}$ is the subgraph of $G$ induced by $V_{i}$ that is either a clique or an independent set, for each $i=1,\ldots,{\tt{nd}}$.
2 The MBV problem
In this section we give an algorithm parameterized by mw  that finds a spanning tree of graph ${{G}}$ with $b({{G}})$ branch vertices. Namely, we prove the following result.
Theorem 1.
Minimum Branch Vertices parameterized by modular-width is fixed-parameter tractable.
Let ${{G}}$ be the input graph. Consider the parse-tree of an algebraic expression describing ${{G}}$, according to the rules (O1)-(O4) in Section 1.3. We take a look at the operation corresponding to the root: Operation (O1) is trivial and (O2) yields a disconnected graph, therefore we suppose the last operation is either (O3) or (O4). Hence, we can see the input graph as ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ where ${{H}}$ is a connected graph with $2\leq{{n}}\leq{\tt{mw}}({{G}})$ vertices and $G_{1},\ldots,G_{{n}}$ are graphs.
The algorithm that finds a spanning tree of an input graph ${{G}}$ with $b({{G}})$ branch vertices goes through the following steps 1) and 2).
  • 1)
    An FPT algorithm for PSC and PP parameterized by the modular-width of the input graph ${{G}}$. Namely, for each vertex ${\hat{{{G}}}}={\hat{{{H}}}}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}})$ of the parse-tree of ${{G}}$, we show how to compute the triple
    $(ham({\hat{{{G}}}}),spi({\hat{{{G}}}}),|V({\hat{{{G}}}})|)$ together with ${\cal{P}}_{ham({\hat{{{G}}}})}$ and ${\cal{P}}_{spi({\hat{{{G}}}})}$,
    where ${\cal{P}}_{ham({\hat{{{G}}}})}$ is a partition of ${\hat{{{G}}}}$ into $ham({\hat{{{G}}}})$ paths and ${\cal{P}}_{spi({\hat{{{G}}}})}$ is a path-spider cover of ${\hat{{{G}}}}$ with $spi({\hat{{{G}}}})-1$ paths.
  • 2)
    Compute a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $b({{G}})$ branch vertices by using the values, computed at step 1), for the graphs $G_{1},\ldots,G_{{n}}$, that is,
    $(ham(G_{i}),spi(G_{i}),|V(G_{i})|)$,  ${\cal{P}}_{ham(G_{i})}$, and ${\cal{P}}_{spi(G_{i})}$,
    for $i=1,\ldots,{{n}}$.
The computation in step 1) is presented in Section 3. The computation in step 2), presented in this section, is only done once, i.e., for the root vertex of the parse tree, corresponding to the input graph ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ whenever $(ham(G_{i}),spi(G_{i}),|V(G_{i})|)$,  ${\cal{P}}_{ham(G_{i})}$, and ${\cal{P}}_{spi(G_{i})}$, for $i=1,\ldots,{{n}}$, are known.
We start by giving a characterization of a spanning tree with the minimum number of branch vertices in terms of the modular decomposition of ${{G}}$.
Lemma 1.
Let ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ be a connected graph. There exists a spanning tree of ${{G}}$ with $b({{G}})$ branch vertices that has at most one branch vertex belonging to $G_{i}$ for each $i=1,\ldots,{{n}}$. Hence, $b({{G}})\leq{{n}}\leq{\tt{mw}}({{G}})$.
Proof.
Let $T$ be a spanning tree of ${{G}}$ with $b({{G}})$ branch vertices.
Denote by $B$ the set of vertices of ${{G}}$ that are branch vertices in $T$ and by $N_{T}(v)$ the set of neighbors of $v$ in $T$, for any vertex $v$. Assume that $|V(G_{i})\cap B|\geq 2$ for some $1\leq i\leq n$. We show how to transform $T$ so to satisfy the lemma. The transformation consists of two phases.
  • Phase 1. For each $i=1,\ldots,n$, we denote by $B_{i}$ the set of branch vertices in $V(G_{i})$ that have in $T$ at least two neighbors outside $V(G_{i})$, that is,
    $B_{i}=\{v\ |\ v\in V(G_{i})\cap B\mbox{ and }|N_{T}(v)\cap(\cup_{j\neq i}V(G_{j}))|\geq 2\}.$
    In this phase we transform $T$ so that $|B_{i}|\leq 1$, for each $i$. We proceed as follows.
    For each $i$ such that $|B_{i}|\geq 2,$
    • let $v$ be any node in $B_{i}$;
    • for each $w\in B_{i}$ with $w\neq v$, consider the path connecting $v$ and $w$ in $T$, say $v,\ldots,w^{\prime},w$, and modify $T$ as follows: For any $x\in(N_{T}(w)\setminus V(G_{i}))\setminus\{w^{\prime}\}$, substitute in $T$ the edge $\{w,x\}$ by the edge $\{v,x\}$.
      (Notice that this is possible by (1) and implies $B_{i}=\{v\}$).
  • Phase 2. We know that each $G_{i}$ contains at most one branch vertex with at least two neighbors outside $V(G_{i})$, that is now $|B_{i}|\leq 1$ for each $i$.
    Suppose there exists $G_{i}$ such that $|V(G_{i})\cap B|\geq 2$. We modify the spanning tree so that only one branch vertex remains among the vertices of $G_{i}$. We proceed as follows.
    While there exists $i$ such that $|V(G_{i})\cap B|\geq 2$.
    • a)
      Choose any $j\neq i$ such that $\{i,j\}\in E(G)$ and let
      $w\in\begin{cases}B_{j}&{\mbox{if $B_{j}=\{w\}$,}}\\ V(G_{j})\cap B&{\mbox{if $B_{j}=\emptyset$ and $V(G_{j})\cap B\neq\emptyset$}}\\ {V(G_{j})}&{\mbox{otherwise.}}\end{cases}$
    • b)
      For each $v\in V(G_{i})\cap B$ with $v\notin B_{i}$, perform the following step.
      • Consider the path connecting $v$ and $w$ in $T$, say $v,v^{\prime},\ldots w$, and modify $T$ as follows: For any $x\in N_{T}(v)\cap V(G_{i})$ and $x\neq v^{\prime}$, substitute in $T$ the edge $\{v,x\}$ by the edge $\{w,x\}$.
    Note that the aforementioned step is facilitated by (1). It’s important to emphasize that even if $w$ becomes a new branch vertex, we observe that $|V(G_{j})\cap B|=1$. Furthermore, $|V(G_{i})\cap B|\leq 1$, ensuring that the total count of branch vertices remains unchanged. Consequently, we have successfully derived a new spanning tree $T^{\prime}$ with a set of branch vertices $B^{\prime}$ such that $|B^{\prime}|\leq|B|$. By repeating steps a) and b), one can construct the desired spanning tree of ${{G}}$ with at most one branch vertex in each $V(G_{i})$. ∎
In the remaining part of this section, we present an algorithm that computes a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $b({{G}})$ branch vertices, if $b({{G}})>0$. In Section 3.2 we deal with the case $b({{G}})=0$, that is, we show how to find a Hamiltonian path of ${{G}}$, if one exists. The work in [18] focuses on verifying the existence of a Hamiltonian path in a graph; the proofs therein implicitly give a method for constructing such a path. We recall such a construction and show its relationship as a special case of our algorithm.
Using Lemma 1, the algorithm proceeds by examining all subsets ${B_{{H}}}\subseteq\{1,\ldots,n\}$ with $|{B_{{H}}}|\geq 1$, ordered by size. For each subset, it checks whether there exists a spanning tree of ${{G}}$ with $|{B_{{H}}}|$ branch vertices such that exactly one branch vertex belongs to each $V(G_{i})$ with $i\in{B_{{H}}}$ and none belongs to $V(G_{i})$ for $i\not\in{B_{{H}}}$.
The identification of the spanning tree of ${{G}}$ involves solving an Integer Linear Program (ILP) that utilizes the values $ham(G_{i}),spi(G_{i}),|V(G_{i})|$, for $i=1,\ldots,{{n}}$, and leverages property (1). Namely, if the ILP does not admit a solution for a subset ${B_{{H}}}$, that subset is discarded. If a solution is found, we demonstrate how to use it along with the partitions of $G_{i}$ given by ${\cal{P}}_{ham(G_{i})}$ and ${\cal{P}}_{spi(G_{i})}$ to construct a spanning tree of ${{G}}$ with exactly $|{B_{{H}}}|$ branch vertices (noting that the subsets ${B_{{H}}}$ are considered in increasing order of size). The optimal spanning tree corresponds to the smallest subset ${B_{{H}}}$ for which the ILP provides a solution.
2.1 The Integer Linear Program
Let ${B_{{H}}}\subseteq\{1,\ldots,{{n}}\}$, with $|{B_{{H}}}|\geq 1$. Construct a digraph
${{{H}}_{B_{{H}}}}=(\{1,\ldots,{{n}}\}\cup\{s\},{A_{B_{{H}}}}),$
where $s\not\in\{1,\ldots,{{n}}\}$ is an additional vertex that will be called the source. ${{{H}}_{B_{{H}}}}$ is obtained from ${{H}}$ by replacing each edge $\{i,j\}\in E({{H}})$ by the two directed arcs $(i,j)$ and $(j,i)$, and then adding a directed arc $(s,r)$ where $r$ is an arbitrary vertex in ${B_{{H}}}$. Formally,
${A_{B_{{H}}}}=\{(s,r)\}\cup\{(i,j),(j,i)\ |\ \mbox{there exists an edge between $i$ and $j$ in }E({{H}})\}.$
For sake of clearness, we will refer to the vertices of ${{H}}$ as module indices and reserve the term vertex to those in ${{G}}$.
We use the solution of the following Integer Linear Programming (ILP) to select arcs of ${{{H}}_{B_{{H}}}}$ that will help to construct the desired spanning tree in ${{G}}$.
$x_{sr}=1$ (2)
$\sum_{j:(j,i)\in{A_{B_{{H}}}}}x_{ji}\leq|V(G_{i})|\qquad\qquad\qquad\forall i\in\{1,\ldots,{{n}}\}$ (3)
$\sum_{j:(j,i)\in{A_{B_{{H}}}}}x_{ji}\geq spi(G_{i})\qquad\qquad\qquad\forall i\in{B_{{H}}}$ (4)
$\sum_{j:(j,i)\in{A_{B_{{H}}}}}x_{ji}\geq ham(G_{i})\qquad\qquad\qquad\forall i\in\{1,\ldots,{{n}}\}\setminus{B_{{H}}}$ (5)
$\sum_{\ell:(i,\ell)\in{A_{B_{{H}}}}}x_{i\ell}-\sum_{j:(j,i)\in{A_{B_{{H}}}}}x_{ji}\leq 0\qquad\ \forall i\in\{1,\ldots,{{n}}\}\setminus{B_{{H}}}$ (6)
$y_{sr}={{n}}$ (7)
$\sum_{j:(j,i)\in{A_{B_{{H}}}}}y_{ji}-\sum_{\ell:(i,\ell)\in{A_{B_{{H}}}}}y_{i\ell}=1\ \qquad\forall i\in\{1,\ldots,{{n}}\}$ (8)
$y_{ij}\leq{{n}}\ x_{ij}\qquad\qquad\qquad\ \ \ \ \ \forall(i,j)\in{A_{B_{{H}}}}$ (9)
$y_{ij},x_{ij}\in\mathbb{N}\qquad\qquad\qquad\ \ \ \ \ \ \forall(i,j)\in{A_{B_{{H}}}}$ (10)
For each arc $(i,j)\in{A_{B_{{H}}}}$, the non-negative decision variable $x_{ij}$ represents the load assigned to $(i,j)$. The load of the arc $(s,r)$ is set to 1. The total incoming load at module index $i\in\{1,\ldots,{{n}}\}$ must be at least $spi(G_{i})$ if $i\in{B_{{H}}}$ (ensuring the spider and all $spi(G_{i})-1$ paths in $G_{i}$ are reached) and at least $ham(G_{i})$ $i\not\in{B_{{H}}}$ (ensuring all $ham(G_{i})$ paths in $G_{i}$ are reached), and it must not exceed $|V(G_{i})|$. Constraints (3), (4) and (5) enforce these requirements.
Constraint (6) ensures that for any $i\not\in{B_{{H}}}$, the outgoing load is upper-bounded by the incoming load.
Constraints (7) and (8) implement a single commodity flow, with $s$ as the source and the other types as demand vertices. For each arc $(i,j)\in{A_{B_{{H}}}}$, the non-negative decision variable $y_{ij}$ represents the flow from $i$ to $j$.
Each $i\in\{1,\ldots,{{n}}\}$ has demand of one unit, meaning the difference between inflow and outflow must be exactly one. Meanwhile, the supply quantity at the source $s$ must be exactly $n$, to reach each of the module index in $\{1,\ldots,{{n}}\}$.
Constraint (9) ensures that $y_{ij}=0$ whenever $x_{ij}=0$; thus, if no load is assigned to $(i,j)$, then $j$ cannot be reached trough $i$.
Given an integer solution $(y,x)$, if any, to the above ILP, the values of variables $y$ ensure that each module index $i\in\{1,\ldots,{{n}}\}$ is reachable from the source $s$. Consequently, by the construction of the digraph ${{{H}}_{B_{{H}}}}$, each module index $i\in\{1,\ldots,{{n}}\}$ is also reachable from module index $r$. Furthermore, due to the relationship between variables $x$ and $y$ (constraint (9)), we know that each module index $i\in\{1,\ldots,{{n}}\}$ receives incoming load from at least one of its neighbors.
Claim 1.
The subgraph ${{{H}}_{x}}$ of ${{{H}}_{B_{{H}}}}$ with vertex set $\{1,\ldots,{{n}}\}$ and arc set $\{(i,j)|$ $x_{ij}\geq 1\}$ contains a directed path from $r$ to any other module index .
We emphasize that the constraints involving variables $y$ only guarantee the existence of a spanning tree in ${{{H}}_{x}}$. A more sophisticated approach is needed to find a spanning tree of ${{G}}$ that has exactly one branch vertex in each $V_{i}$ with $i\in{B_{{H}}}$.
2.2 The spanning tree construction
Our algorithm constructs a spanning tree $T$ of ${{G}}$ with $|{B_{{H}}}|$ branch vertices, one in each $V(G_{i})$ with $i\in{B_{{H}}}$. To this aim, it uses the values of variables $x$ and the path-spider cover ${\cal{P}}_{spi(G_{i})}$ for $i\in{B_{{H}}}$ and the partition into disjoint paths ${\cal{P}}_{ham(G_{i})}$ for $i\not\in{B_{{H}}}$.
Denote by $In(i)$ the set of the module indices for which there exist arcs in ${{{H}}_{x}}$ toward $i$, that is, $In(i)=\{j\ |x_{ji}\geq 1\}$, and by
${\alpha}_{i}=\sum_{j:j\in In(i)}x_{ji}$ (11)
the number of vertices of $V(G_{i})$ whose parent in $T$ is a vertex outside $V(G_{i})$.
Let ${\cal{P}}^{i}=\{P_{1}^{i},P_{2}^{i},\ldots,P_{{\alpha}_{i}}^{i}\}$ be
  • $\bullet$
    the path-spider cover of $G_{i}$ obtained from those in ${\cal{P}}_{spi(G_{i})}$ by removing ${\alpha}_{i}-spi(G_{i})$ arbitrary edges in case $i\in{B_{{H}}}$ (notice that by constraint (4), it holds ${\alpha}_{i}\geq spi(G_{i})$), or
  • $\bullet$
    the partition of $G_{i}$ into disjoint paths obtained from those in ${\cal{P}}_{ham(G_{i})}$ by removing ${\alpha}_{i}-ham(G_{i})$ arbitrary edges in case $i\not\in{B_{{H}}}$ (notice that by (5), it holds ${\alpha}_{i}\geq ham(G_{i})$).
Furthermore, denote by
$f({\cal{P}}^{i})=\{f(P_{1}^{i}),f(P_{2}^{i}),\ldots,f(P_{{\alpha}_{i}}^{i})\}$
the sets of the first end-points in the partition ${\cal{P}}^{i}$ and by
$s({\cal{P}}^{i})=\{s(P_{1}^{i}),s(P_{2}^{i}),\ldots,s(P_{{\alpha}_{i}}^{i})\}$
the sets of the second end-points in ${\cal{P}}^{i}$. In case $i\in{B_{{H}}}$, we assume that $P_{1}^{i}\in{\cal{P}}^{i}$ is the spider and $f(P_{1}^{i})=s(P_{1}^{i})$ is the center in $P_{1}^{i}$.
We also denote by
$\beta_{i}=\begin{cases}\sum_{\ell:i\in In(\ell)}x_{i\ell}&\text{if $i\not\in{B_{{H}}}$}\\ 1&\text{if $i\in{B_{{H}}}$}\end{cases}$ (12)
the number of vertices of $V(G_{i})$, that will be the parent of some vertex in $\bigcup_{\ell:i\in In(\ell)}V(G_{\ell})$.
Our algorithm ensures that the ${\alpha}_{i}$ vertices in $f({\cal{P}}^{i})$ are the vertices in $G_{i}$ whose parent in $T$ is located outside $V(G_{i})$, and that $\beta_{i}$ vertices from $s({\cal{P}}^{i})$ are selected to be parents of vertices outside $V(G_{i})$. Notice that by Claim 1 (${\alpha}_{i}\geq 1$) and constraint (6), it follows that ${\alpha}_{i}\geq\beta_{i}$ for each $i=1,\ldots,{{n}}$.
Figure 3 shows the partition of graph $G_{i}$ (whose vertices are grouped in the dotted circle) into ${\alpha}_{i}$ disjoint paths if $i\not\in{B_{{H}}}$ and into a spider plus ${\alpha}_{i}-1$ disjoint paths if $i\in{B_{{H}}}$.
Figure 3: The vertices of graph $G_{i}$, grouped in the dotted circle, as partitioned in ${\alpha}_{i}$ disjoint paths if $i\not\in{B_{{H}}}$ and in a spider plus ${\alpha}_{i}-1$ disjoint paths if $i\in{B_{{H}}}$. Vertex $f(P_{j}^{i})$ is the only vertex in $P_{j}^{i}$ whose parent in $T$ is outside $G_{i}$ and vertex $s(P_{j}^{i})$ is the only vertex in $P_{j}^{i}$ that can have a children in $T$ outside $G_{i}$.
The TREE algorithm iteratively constructs a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ by exploring unexplored vertices of ${{G}}$, until all vertices are explored. It maintains a main subtree $T$ and a forest, with the roots of these trees progressively connected to $T$ to form the spanning tree. The process halts when all vertices of ${{G}}$ have been explored.
The algorithm maintains a set $R$, which contains the roots of trees with explored vertices that are waiting to be connected to the main tree $T$. The structure of the forest is described using the parent function $\pi$. At the beginning the set $R$ is empty. The exploration begins with the center $f(P_{1}^{r})$ of the spider in the path-spider cover of $G_{r}$ (recall that $r\in{B_{{H}}}$ by construction ). The procedure EXPLORE$(f(P_{1}^{r}))$ constructs the main tree $T$ rooted at $f(P_{1}^{r})$ and marks all reached vertices as explored, adding them to the set $Ex$. Clearly, for each explored vertex $v$, there is a path in $T$ that joins $f(P_{1}^{r})$ to $v$. The algorithm EXPLORE uses a queue $Q$ to enqueue the explored vertices.
However, it is possible that some vertices remain unexplored (i.e., $V({{G}})\setminus Ex\neq\emptyset$). In such scenarios, there exists a module $G_{j}$ where an explored vertex $w\in f({\cal{P}}^{j})\cap Ex$ exists alongside an unexplored vertex $u\in(f({\cal{P}}^{j})\setminus Ex)\setminus R$, capable of exploring at least one unexplored neighbor outside $G_{j}$, that is, $\beta_{j}\geq 1$ (the existence of such a set $V(G_{j})$ is assured by Lemma 4). Utilizing (1) and leveraging the knowledge that parents of vertices in $f({\cal{P}}^{j})$ are outside $V(G_{j})$, the algorithm makes:
  • the parent of $w$ (recall that $w$ is explored) becomes the parent of $u$, and
  • $w$, the root of a subtree containing explored vertices, is added to $R$ and removed from $Ex$. This action enables $w$ to be later explored and added, along with its subtree, to the main tree $T$.
  • EXPLORE$(u)$ is invoked to initiate a new exploration starting from $u$.
Notice that the algorithm updates the forest by assigning the parent of $w$ to $u$. Only later, after adding $u$ and some of its descendants, the subtree rooted at $w$ is reintegrated into the main tree $T$. This approach allows connecting new vertices in $V({{G}})\setminus Ex$ to the main tree $T$. The specific selection of $u$ and $w$ is designed to prevent scenarios where the algorithm could fail, ensuring that no arc can be added to $T$ without either forming a cycle or introducing an additional branch vertex. The process continues iteratively as long as there are unexplored vertices, i.e., $V({{G}})\setminus Ex\neq\emptyset$.
The EXPLORE$(u)$ procedure initiates exploration from $u=f(P^{j}_{k})$ along with the entire path or spider $P^{j}_{k}$ Assume that when in the algorithm $P\in{\cal{P}}^{j}$ is explored (i.e., $Ex=Ex\cup P$), the parent function $\pi$ is updated. For path, $\pi$ is set from $f(P)$ to $s(P)$, and for spiders, it is set from the center $f(P)=s(P)$ to the leaves.. Initially, only the vertex $s(P^{j}_{k})$ is placed in $Q$. Subsequently, the main tree $T$ is expanded. If $u\neq f(P^{r}_{1})$, the parent of $u$ is already a vertex in $T$, facilitating the construction of a subtree rooted at $u$ that spans all newly explored vertices.
EXPLORE$(u)$ utilizes the values of $\alpha_{i}$ and $\beta_{i}$ for each module index $i$, defined initially as in (11) and (12), and the partition ${\cal{P}}^{i}=\{P_{1}^{i},P_{2}^{i},\ldots,P_{{\alpha}_{i}}^{i}\}$ of $G_{i}$.
The value ${\alpha}_{i}=\sum_{j:j\in In(i)}x_{ji}$ counts the number of vertices of $V(G_{i})$ that must have a parent outside $V(G_{i})$, specifically $f({\cal{P}}^{i})=\{f(P_{1}^{i}),f(P_{2}^{i}),\ldots,f(P_{{\alpha}_{i}}^{i})\}$. In particular, $x_{ji}$ vertices from $f({\cal{P}}^{i})$ need to be explored by vertices in $V(G_{j})$, for $j\in In(i)$. The value $\beta_{i}$ counts the number of vertices of $V(G_{i})$ that need to explore other vertices in some other $V(G_{\ell})$, for $\ell:i\in In(\ell)$. In particular,
    if $i\in{B_{{H}}}$ then exactly $\beta_{i}=1$ vertex in $V(G_{i})$, that is $s(P^{i}_{1})$ (i.e., the center of the spider $P^{i}_{1}$), becomes a branch vertex in $T$: it is set as the parent of $x_{i\ell}$ unexplored vertices in $f({\cal{P}}^{\ell})$ for each $\ell$ such that $x_{i\ell}\geq 1$ (i.e., $i\in In(\ell)$), and
    if $i\not\in{B_{{H}}}$ then $\beta_{i}=\sum_{\ell:i\in In(\ell)}x_{i\ell}$ vertices in $s({\cal{P}}^{i})=\{s(P_{1}^{i}),s(P_{2}^{i}),\ldots,s(P_{{\alpha}_{i}}^{i})\}$ are chosen and each one becomes the parent of one unexplored vertex in $f({\cal{P}}^{\ell})$.
    Recall that, by the ILP constraints, we know that $\alpha_{i}\geq\beta_{i}$.
Only vertices in $s({\cal{P}}^{i})$ for $1\leq i\leq{{n}}$ are enqueued in $Q$. When a vertex $v\in s({\cal{P}}^{i})$ is dequeued from $Q$ in EXPLORE$(u)$ then the value of $\beta_{i}$ is decreased by one if $v$ explores (i.e., if $\beta_{i}\geq 1$). In such cases, for each explored vertex $f(P^{\ell}_{h})$ where $i\in In(\ell)$, the entire path $P^{\ell}_{h}$ is also explored.
Additionally, the value $\alpha_{\ell}$ is reduced by the number of vertices in $s({\cal{P}}^{\ell})$ that $v$ explores, for $i\in In(\ell)$. Therefore, at the start of each iteration of the while loop in EXPLORE$(u)$ the value $\alpha_{i}$ represents the number of vertices in $f({\cal{P}}^{i})$ that remain to be explored while $\beta_{i}$ indicates the number of vertices in $s({\cal{P}}^{i})$ that already need to explore. It’s noteworthy that when a vertex $v\in s({\cal{P}}^{i})$ is dequeued from $Q$ in EXPLORE$(u)$, with $u\neq f(P^{r}_{1})$, and $\beta_{i}\geq 1$, the algorithm verifies if the neighbour $v^{\prime}\in f({\cal{P}}^{\ell})$, which $v$ explores, is in $R$ (i.e., $v^{\prime}$ is a root of a tree in the forest). If so, $v^{\prime}$ and its subtree are connected to the main tree $T$, as it has already been explored in previous iterations.
Algorithm 1 TREE$({{G}},G_{1},\ldots,G_{{n}},r,{B_{{H}}})$
1:$R=\emptyset$,   $B=\emptyset$,   $Ex=\emptyset$
2:$\pi(u)=nhil$   for each $u\in V({{G}})$
3:EXPLORE$(f(P_{1}^{r}))$
4:while $V({{G}})\setminus Ex\neq\emptyset$ do
5:     - Let $G_{j}$ be any graph s.t. (($f({\cal{P}}^{j})\setminus Ex)\setminus R\neq\emptyset\neq f({\cal{P}}^{j})\cap Ex$) and $\beta_{j}\geq 1$
6:     - Let $w\in f({\cal{P}}^{j})\cap Ex$ and $u\in(f({\cal{P}}^{j})\setminus Ex)\setminus R$
7:     - Set $\pi(u)=\pi(w)$,    $Ex=Ex-\{w\}$,    $R=R\cup\{w\}$
8:     - EXPLORE$(u)$
9:end while
10:return $\pi$,   $B$
Algorithm 2 EXPLORE$(u)$
1:Let $Q$ be an empty queue
2:Let $u=f(P^{j}_{k})$
3:$Ex=Ex\cup P^{j}_{k}$
4:$Q.enqueue(s(P^{j}_{k}))$
5:while $Q\neq\emptyset$ do
6:     $v=Q.dequeue$
7:     Let $v\in s({\cal{P}}^{i})$
8:     if $i\notin{B_{{H}}}$ and $\beta_{i}\geq 1$  then
9:         - Let $f(P^{\ell}_{h})\in f({\cal{P}}^{\ell})\setminus Ex$ for some $\ell$ s.t. $i\in In(\ell)$
10:         - $\pi(f(P^{\ell}_{h}))=v$
11:         if $f(P^{\ell}_{h})\not\in R$ then
12:              - $Ex=Ex\cup P^{\ell}_{h}$
13:              - $Q.enqueue(s(P^{\ell}_{h}))$
14:         else  $R=R\setminus\{f(P^{\ell}_{h})\}$
15:         end if
16:         - ${\alpha}_{\ell}={\alpha}_{\ell}-1$
17:         - $\beta_{i}=\beta_{i}-1$
18:     else if $i\in{B_{{H}}}$ and $\beta_{i}=1$ then
19:         - $B=B\cup\{v\}$
20:         for each $\ell$ s.t. $i\in In(\ell)$ do
21:              - Let $A_{i\ell}\subseteq f({\cal{P}}^{\ell})-Ex$ s.t. $|A_{i\ell}|=x_{i\ell}$
22:              - ${\alpha}_{\ell}={\alpha}_{\ell}-x_{i\ell}$,
23:              for each $f(P^{\ell}_{h})\in A_{i\ell}$ do
24:                  - $\pi(f(P^{\ell}_{h}))=v$
25:                  if $f(P^{\ell}_{h})\not\in R$ then
26:                       -$Ex=Ex\cup P^{\ell}_{h}$
27:                       -$Q.enqueue(s(P^{\ell}_{h}))$
28:                  else  $R=R\setminus\{f(P^{\ell}_{h})\}$
29:                  end if
30:              end for
31:         end for
32:         - $\beta_{i}=\beta_{i}-1$
33:     end if
34:end while
Lemma 2.
At the end of EXPLORE$(f(P^{r}_{1}))$ the function $\pi$ describes a tree, rooted at $f(P^{r}_{1})$, spanning the set $Ex\subseteq V(H)$ of explored vertices. The vertices in $B\cap Ex$ are the branch vertices.
Proof.
When EXPLORE$(f(P^{r}_{1}))$ is is invoked, the entire spider $P^{r}_{1}$ is explored (i.e. $Ex=Ex\cup P^{r}_{1}$ and it is added to $T$) and its center $s(P^{r}_{1})$ is enqueued in $Q$. Subsequently, each time a vertex $v\in s({\cal{P}}^{i})$ is dequeued from $Q$ (recall, $v\in Ex$, indicating it is an explored vertex), the algorithm can either terminate its exploration (i.e., $\beta_{i}=0$) or explore unexplored neighbors of $v$ together with the path/spider it belongs to. It can be shown that $v$ indeed has the necessary number of unexplored neighbors. If $\beta_{i}=0$ then $v$ is a leaf in $T$, and only the case where $\beta_{i}\geq 1$ needs consideration. If $i\not\in{B_{{H}}}$ then $v$ has $\beta_{i}\geq 1$ unexplored neighbors and one of them, say $f(P^{\ell}_{h})$ for $i\in In(\ell)$, can be added to $T$ as child of $v$ If $i\in{B_{{H}}}$ then $v$ is the first vertex of $V(G_{i})$ to explore, and $x_{i\ell}$ vertices in $f({\cal{P}}_{\ell})$ are unexplored and can be added to $T$ as children of $v$, for each $\ell$ such that $x_{i\ell}\geq 1$. Therefore, $v$ becomes a branch vertex in $T$ and is added to $B$. Since $R=\emptyset$ (indicating no trees are in the forest), each time a neighbor of $v$ is explored, such as $f(P^{\ell}_{h})$, the entire path/spider $P^{\ell}_{h}$ is added to $T$ and $s(P^{\ell}_{h})$ is enqueued in $Q$. Then, every explored vertex has $f(P^{r}_{1})$ as its ancestor, meaning the function $\pi$ defines a path form any explored vertex to $f(P^{r}_{1})$. Since no vertex can be enqueued in $Q$ more than once (as each enqueued vertex is also marked as explored), the function $\pi$ ensures that no cycles are created. ∎
Fix any iteration of the while loop in algorithm TREE and define ${{{H}}^{\prime}_{x}}$ as the subgraph of ${{{H}}_{x}}$ (refer to Claim 1) containing the arc $(i,j)$ if, at the beginning of the while loop, fewer than $x_{ij}$ vertices in $f({\cal{P}}^{j})$ have been assigned a parent in $s({\cal{P}}^{i})$.
Lemma 3.
Let $j$ be any module index in ${{{H}}^{\prime}_{x}}$. The following properties hold for ${{{H}}^{\prime}_{x}}$:
  • (a)
    If $f({\cal{P}}^{j})\subseteq Ex$ then the module index $j$ is isolated in ${{{H}}^{\prime}_{x}}$.
  • (b)
    The module index $r$ s.t. $f(P^{r})$ contains the root $f(P^{r}_{1})$ of $T$ has no outgoing arcs in ${{{H}}^{\prime}_{x}}$.
  • (c)
    $f({\cal{P}}^{j})\not\subseteq Ex$ if and only if $j$ has at least one incoming arc in ${{{H}}^{\prime}_{x}}$.
  • (d)
    If $f({\cal{P}}^{j})\cap Ex=\emptyset$ then $j$ keeps in ${{{H}}^{\prime}_{x}}$ all the incoming and outgoing arcs it has in ${{{H}}_{x}}$.
  • (e)
    If $f({\cal{P}}^{j})\setminus Ex\subseteq R$ then $j$ has no outgoing arcs in ${{{H}}^{\prime}_{x}}$.
Proof.
Consider property (a). If each vertex in $f({\cal{P}}^{j})$ is explored, then it has a parent in $T$. Recalling that $\sum_{t:t\in In(j)}x_{tj}={\alpha}_{j}=|f({\cal{P}}^{j})|$, it follows that $j$ has no incoming arc in ${{{H}}^{\prime}_{x}}$. Moreover, procedure EXPLORE assures that once a vertex in $f({\cal{P}}^{j})$ is explored (i.e. it has assigned a parent) then a vertex in $s({\cal{P}}^{j})$ will be assigned at least one child as long as there are vertices to be explored from $V(G_{j})$ (i.e., if $\beta_{j}\geq 1$). In particular, if $j\not\in{B_{{H}}}$ then since $\sum_{t:j\in In(t)}x_{jt}=\beta_{j}=|s({\cal{P}}^{j})|$, it follows that $x_{jt}$ vertices in $f({\cal{P}}^{t})$ have a parent in $s({\cal{P}}^{j})$, for each $t$ such that $j\in In(t)$. Hence $j$ has no outgoing arc surviving in ${{{H}}^{\prime}_{x}}$. If, otherwise, $j\in{B_{{H}}}$ then vertex $s(P^{j}_{1})$ has $x_{jt}>0$ children in $f({\cal{P}}^{t})$ for each $t$ such that $j\in In(t)$. Hence, also in this case $j$ has no outgoing arc surviving in ${{{H}}^{\prime}_{x}}$.
Property (b) follows from the observation that, by construction, $f(P^{r}_{1})$ is a branch vertex of $T$ and has $x_{rt}$ children in each $f({\cal{P}}^{t})$ such that $x_{rt}>0$, (i.e., $r\in In(t)$).
Property (c) follows from the observation that $f({\cal{P}}^{j})\not\subseteq Ex$ is equivalent to say that
$\sum_{t:t\in In(j)}x_{tj}={\alpha}_{j}=|f({\cal{P}}^{j})|>|f({\cal{P}}^{j})\cap Ex|.$
Property (d) follows from the observation that $f({\cal{P}}^{j})\cap Ex=\emptyset$ implies that $j$ still has an incoming neighbour for each $t$ such that $x_{tj}>0$ and an outgoing neighbor for each $t$ such that $x_{jt}>0$.
Property (e) follows from the observation that when the algorithm TREE disconnects a vertex $w$ and adds it to $R$, vertex $w$ has already been assigned its child/children. Hence, if $f({\cal{P}}^{j})$ does not contain any unexplored vertex outside $R$ then $\beta_{j}$ has been decreased to 0. This means that all the $x_{tj}$ arcs from a vertex in $s({\cal{P}}^{t})$ to one in $f({\cal{P}}^{j})$ have been added to the forest, for each $j=1,\ldots,{{n}}$. ∎
We can prove the following results.
Lemma 4.
Let $Ex$ be the set of explored vertices at the beginning of any iteration of the while loop in algorithm TREE. If $V({{G}})\setminus Ex\neq\emptyset$ then there exists a module index $j$ such that $(f({\cal{P}}^{j})\setminus Ex)\setminus R\neq\emptyset\neq f({\cal{P}}^{j})\cap Ex$ and $\beta_{j}\geq 1$.
Proof.
By (a) and (c) of Lemma 3, we know that each module index $j$ in ${{{H}}^{\prime}_{x}}$ is either isolated or has at least one incoming arc. Hence, we focus on the subset of non-isolated module indices . Since each of them has an incoming arc, ${{{H}}^{\prime}_{x}}$ contains a cycle. Consequently, each module index $j$ on such a cycle has an outgoing arc and satisfies $\beta_{j}\geq 1$. Furthermore, by (e) of Lemma 3, each module index $j$ on such a cycle satisfies $(f({\cal{P}}^{j})\setminus Ex)\setminus R\neq\emptyset$.
We show now that at least one module index $j$ on the cycle has $f({\cal{P}}^{j})\cap Ex\neq\emptyset$. Point (b) of Lemma 3 implies that ${{{H}}^{\prime}_{x}}$ does not contain any path from $r$ to any module index on the cycle. If we suppose that for each module index $j$ in the cycle $f({\cal{P}}^{j})\cap Ex=\emptyset$, then (d) of Lemma 3 implies that also ${{{H}}_{x}}$ does not contain a path from $r$ to $j$, thus contradicting Claim 1. ∎
Lemma 5.
After each call of EXPLORE$(u)$ the function $\pi$ describes a forest spanning the vertices in $Ex\cup R$ of explored vertices and consisting of $|R|+1$ trees respectively rooted at $f(P^{r}_{1})$ and at the vertices in $R$. The vertices in $B$ are the only branch vertices in the forest.
Proof.
When EXPLORE$(u)$ is called, the function $\pi$ describes a forest, spanning the current set $Ex\cup R$ with roots in $\{f(P^{r}_{1})\}\cup R$, where $R\subset V({{G}})\setminus Ex$. According to Lemma 2, this is true the first time EXPLORE is called, that is, after the call to EXPLORE$(f(P^{r}_{1}))$ (at that time $R=\emptyset$).
We will prove that the claim holds at the end of each call to EXPLORE$(u)$. When EXPLORE$(u)$ is invoked, $Q$ is initially empty. The vertex $u$ is explored (i.e. added to $Ex$) and enqueued in $Q$. Then EXPLORE$(u)$ proceeds similarly to EXPLORE$(f(P^{r}_{1}))$, dequeueing vertices from $Q$ and exploring their unexplored neighbors, thereby constructing a subtree of the main tree $T$ rooted at $u$ described by the function $\pi$. The only difference from EXPLORE$(f(P^{r}_{1}))$ occurs when one of the explored vertices is $v^{\prime}\in R$. Vertex $v^{\prime}\in R$ is removed from $R$ (see lines 14, 28), connected to the main tree $T$ through the function $\pi$, and marked as explored like any other explored vertex. However $v^{\prime}$ is not enqueued in $Q$ because it has already explored its neighbors; thus, $v^{\prime}$ is connected to $T$ along with its subtree of explored vertices. ∎
Lemma 6.
The algorithm TREE returns a spanning tree of ${{G}}$, described by function $\pi$, with branch vertex set $B$.
Proof.
By Lemma 2 we know that algorithm TREE constructs a main tree $T$ through the procedure EXPLORE$(f(P^{r}_{1}))$, which is described by $\pi$. If $T$ does not span all the vertices in $V({{G}})$ then, Lemma 4 assures that the algorithm finds a graph $G_{j}$ with an explored vertex $w\in f({\cal{P}}^{j})\cap Ex$ and an unexplored vertex $u\in(f({\cal{P}}^{j})\setminus Ex)\setminus R$. By disconnecting $w$ (along its subtree) from the main tree $T$, the algorithm allows $w$ to become one of the roots of trees in $R$. Additionally, since the parent of $w$ in $T$ is a vertex outside $V(G_{j})$ and, since $u$ and $w$ share the same neighborhood outside $G_{j}$ (as indicated by (1)), the algorithm connects $u$ to the vertex that was the parent of $w$ in $T$ (thereby connecting $u$ to $T$). Since $u\not\in R$ and $\beta_{j}\geq 1$, the algorithm starts a new exploration from $u$ (recall that $u\in f({\cal{P}}^{j})\setminus Ex$) by calling EXPLORE$(u)$. By Lemma 5, this allows $T$ to be padded with the subtree rooted a $u$. The lemma follows by iterating the above procedure until no unexplored vertex exists in $V({{G}})$. ∎
2.3 The algorithm complexity
In summary, given the triple $(ham(G_{i}),spi(G_{i}),|V(G_{i})|)$ along with ${\cal{P}}_{ham(G_{i})}$ and ${\cal{P}}_{spi(G_{i})}$, for each $i=1,\ldots,{{n}}$, the proposed method for constructing the spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ works as follows.
For each ${B_{{H}}}\subseteq\{1,\ldots,{{n}}\}$ with $|{B_{{H}}}|\geq 1$, selected in order of increasing size, the algorithm performs the following steps:
  • solve the corresponding ILP
  • if a solution exists for the current set ${B_{{H}}}$, use algorithm TREE to construct a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $|{B_{{H}}}|$ branch vertices.
Jansen and Rohwedderb [29] have recently showed that the time needed to find a feasible solution of an ILP with $p$ integer variables and $q$ constraints is $O(\sqrt{q}\Delta)^{(1+o(1))q}+O(qp)$, where $\Delta$ is the largest absolute value of any coefficient in the ILP. Denoted by ${{m}}$ the number of edges of ${{H}}$, our ILP has $q=3{{n}}+2{{m}}+1$ constraints, $p=2({{m}}+1)$ variables and $\Delta={{n}}$. Hence the time to solve it is $O({{n}}\sqrt{{{n}}+{{m}}})^{(1+o(1))(3{{n}}+2{{m}}+1)}+O({{n}}({{n}}+{{m}})).$ Using the solution $(y,x)$ of the ILP, the algorithm TREE returns the spanning tree of ${{G}}$ in time $O(|V({{G}})|^{2})$. Overall, the algorithm requires time
$2^{{n}}[O({{n}}\sqrt{{{n}}+{{m}}})^{(1+o(1))(3{{n}}+2{{m}}+1)}+O({{n}}({{n}}+{{m}}))]+O(|V({{G}})|^{2}).$
Recall that $n\leq{\tt{mw}}$, and therefore $m\leq{\tt{mw}}^{2}$.
2.4 Optimality
In this section we prove that the spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with the minimum number of branch vertices is the tree that can be obtained by using the smallest set ${B_{{H}}}\subseteq\{1,\ldots,{{n}}\}$ for which the ILP admits a solution. Namely, we can prove the following result.
Lemma 7.
Given the graphs $G_{1},\ldots,G_{{n}}$ and $ham(G_{i}),spi(G_{i}),|V(G_{i})|$ for each $i=1,\dots,{{n}}$. Let $T$ be the spanning tree in ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $b(G)$ branch vertices. Then there exists a set ${B_{{H}}}\subseteq\{1,\ldots,{{n}}\}$ with $|{B_{{H}}}|=b(G)$, for which ILP admits a solution $(x,y)$.
Proof.
Let $B$ be the set of branch vertices in $T$ and $|B|=b(G)$. We show how to obtain from $T$ and $B$ an assignment of values to the variables in $x$ and $y$ that satisfy the constraint (2)-(9) of ILP.
By Lemma 1 we can assume that $|B\cap V(G_{i})|\leq 1$ for each $i=1,\ldots,{{n}}$. Let ${B_{{H}}}=\{i\ |\ B\cap V(G_{i})\neq\emptyset,\ i=1,\ldots,{{n}}\}$. Choose any $r\in{B_{{H}}}$ and let $u_{r}$ the branch vertex in $V(G_{r})$. Root $T$ at $u_{r}$ and direct each edge in $T$ so that there is a path of directed arcs from $u_{r}$ to any vertex $u\in V({{G}})\setminus\{u_{r}\}$. Let $A_{T}$ be the set of all the arcs in $T$.
We set $x_{sr}=1$ (satisfying constraint (2) of ILP), and for $i,j\in\{1,\ldots,{{n}}\}$,
$x_{ij}=|\{(u,v)\ |\ (u,v)\in A_{T},\ u\in V(G_{i}),\ v\in V(G_{j})\}|.$
Let $In(i)=\{j\ |\ x_{ji}\geq 1\}$, for $i=1,\ldots,{{n}}$. Since each vertex $u\in V(G_{i})$ has a parent in $T$, we have the parent of any $u\in V(G_{i})$ can be either a vertex in $V(G_{i})$ or a vertex in $V(G_{j})$ with $j\in In(i)$. This implies that
$\sum_{j:(j,i)}x_{ji}\leq|V(G_{i})|\qquad\qquad\qquad\mbox{for each $i\in\{1,\ldots{{n}}\}$}.$
satisfying constraint (3) of ILP.
Furthermore, if $V(G_{i})$ does not contain a branch vertex then the tree $T$ induces at most $ham(G_{i})$ disjoint paths in $G_{i}$. Hence, at least $ham(G_{i})$ vertices in $V(G_{i})$ have a parent in some $V(G_{j})$ with $j\in In(i)$.
$\sum_{j:(j,i)}x_{ji}\geq ham(G_{i})\qquad\qquad\qquad\mbox{for each $i\in\{1,\ldots,{{n}}\}\setminus{B_{{H}}}$}$
satisfying constraint (5) of ILP.
While, if $V(G_{i})$ contains a branch vertex then $T$ induces a spider and at most $spi(G_{i})-1$ disjoint paths in $G_{i}$. Hence, at least $spi(G_{i})$ vertices in $V(G_{i})$ have a parent in some $V(G_{j})$ with $j\in In(i)$.
$\sum_{j:(j,i)}x_{ji}\geq spi(G_{i})\qquad\qquad\qquad\mbox{for each $i\in{B_{{H}}}$}$
satisfying constraint (4) of ILP.
If $i\not\in{B_{{H}}}$ then $V(G_{i})$ does not contain branch vertices. Hence, each vertex $u\in V(G_{i})$ can be the parent of at most one vertex. Hence,
$\sum_{\ell:(i,\ell)}x_{i\ell}\leq\sum_{j:(j,i)}x_{ji}\qquad\qquad\qquad\mbox{for each $i\in\{1,\ldots,{{n}}\}\setminus{B_{{H}}}$}$
satisfying constraint (6) of ILP.
To assign values to the variables $y$, we introduce the digraph ${{H}}^{x}$ having vertex set $\{1,\ldots,{{n}}\}$ and arc set $\{(i,j)\ |\ x_{ij}\geq 1\}$. Let $T_{x}$ be the tree rooted at $r$ obtained by a BFS visit of ${{H}}^{x}$. For each $i\in\{1,\ldots,{{n}}\}\setminus\{r\}$, let $p(i)$ the parent of $i$ in $T_{x}$. Pad $T_{x}$, adding arc $(s,r)$ (i.e., $p(r)=s$). We set $y_{sr}={{n}}$ (satisfying constraint (7) of ILP) and for $i\in\{1,\ldots,{{n}}\}\setminus\{r\}$ we set
$y_{ji}=\begin{cases}{\mbox{the number of vertices in the subtree of $T_{x}$ rooted at $i$}}&{\mbox{if $j=p(i)$}}\\ {0}&{\mbox{if $j\neq p(i)$}}\end{cases}$
Hence,
$\sum_{j:(j,i)}y_{ji}=y_{p(i)i}=1+\sum_{\ell:p(\ell)=i}y_{i\ell}=1+\sum_{\ell:(i,\ell)}y_{i\ell}$
satisfying constraint (8) of ILP.
We notice that the number of vertices in the subtree of $T_{x}$ rooted at $i$ is at most ${{n}}$, for each $i\in\{1,\ldots,{{n}}\}$. Moreover, recalling that $x_{p(i)i}\geq 1$, we know that ${{H}}^{x}$ contains $(p(i),i)$. Therefore, we get
$y_{p(i)i}\leq{{n}}\leq{{n}}\ x_{p(i)i}$
satisfying constraint (9) of ILP since $y_{ji}=0$ for each $j\neq p(i)$. ∎
3 The triple and partition computation
Following [18], we use a bottom-up dynamic programming approach along the parse-tree of an algebraic expression describing ${{G}}$, to compute for every internal vertex ${\hat{{{G}}}}={\hat{{{H}}}}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}})$ of the parse tree a record of data, using those already computed for its children. Specifically, given the triple $(ham({\hat{G}}_{i}),spi({\hat{G}}_{i}),|V({\hat{G}}_{i})|)$,  ${\cal{P}}_{ham({\hat{G}}_{i})}$ and ${\cal{P}}_{spi({\hat{G}}_{i})}$, for each $i=1,\ldots,{\hat{n}}$, we need to compute
the triple $(ham({\hat{{{G}}}}),spi({\hat{{{G}}}}),|V({\hat{{{G}}}})|)$ together with ${\cal{P}}_{ham({\hat{{{G}}}})}$ and ${\cal{P}}_{spi({\hat{{{G}}}})}$.
Notice that in case ${\hat{G}}_{i}$, for $i=1,\ldots,{\hat{n}}$, is a leaf in the parse tree (i.e., ${\hat{G}}_{i}=(\{v\},\emptyset)$ for some $v\in V({\hat{{{G}}}})$- operation (O1)) then $ham({\hat{G}}_{i})=spi({\hat{G}}_{i})=1$ and ${\cal{P}}_{ham({\hat{{{G}}}})}={\cal{P}}_{spi({\hat{{{G}}}})}=\{v\}$.
Clearly, $|V({\hat{{{G}}}})|=\sum_{i=1}^{\hat{n}}|V({\hat{G}}_{i})|$. Below, we show how to compute $spi({\hat{{{G}}}})$ and ${\cal{P}}_{spi({\hat{{{G}}}})}$, and also $ham({\hat{{{G}}}})$ and ${\cal{P}}_{ham({\hat{{{G}}}})}$.
[…]
3.1 Computing $spi({\hat{{{G}}}})$ and ${\cal{P}}_{spi({\hat{{{G}}}})}$
In order to compute the values $spi({\hat{{{G}}}})$ and ${\cal{P}}_{spi({\hat{{{G}}}})}$, we first need a preliminary result.
Fact 1.
Let ${\hat{{{G}}}}$ be a graph and
$s({\hat{{{G}}}})=\min\{\ell\ |\ {\hat{{{G}}}}\otimes(\ell-1)\;\mbox{has a spanning spider with center in ${\hat{{{G}}}}$}\}.$
Then $spi({\hat{{{G}}}})=s({\hat{{{G}}}})$.
Proof.
We first show that $s({\hat{{{G}}}})\leq spi({\hat{{{G}}}})$. Consider $P_{1},P_{2},\ldots,P_{spi({\hat{{{G}}}})}$, the path-spider cover of ${\hat{{{G}}}}$, where $f(P_{1})$ is the center of spider $P_{1}$ and, $f(P_{i})$ denotes the first end-point of path $P_{i}$ for $i=2,\ldots,spi({\hat{{{G}}}})$. Consequently, the graph ${\hat{{{G}}}}\otimes(spi({\hat{{{G}}}})-1)$ contains a spider with center $f(P_{1})$, connecting $f(P_{1})$ to the vertex $i$, and then connecting vertex $i$ to $f(P_{i})$ for each for $i=1,\ldots,spi({\hat{{{G}}}})-1$. Next, we establish that $spi({\hat{{{G}}}})\leq s({\hat{{{G}}}})$. Let $S$ be a spider in ${\hat{{{G}}}}\otimes(s({\hat{{{G}}}})-1)$ with center $u\in V({\hat{{{G}}}})$. Removing the vertices in $\{1,\ldots,s({\hat{{{G}}}})-1\}$ from $S$, we obtain a path-spider cover with center $u$ and $s({\hat{{{G}}}})-1$ pairwise disjoint paths. ∎
Recall that the graph ${\hat{{{G}}}}\otimes(\ell-1)$ is equal to ${\hat{{{H}}}}^{\prime}({\hat{G}}_{1},\ldots,{\hat{G}}_{n},I_{\ell-1})$ and notice that
$ham(I_{\ell-1})=spi(I_{\ell-1})=\ell-1\mbox{ and }{\cal{P}}_{ham(I_{\ell-1})}={\cal{P}}_{spi(I_{\ell-1})}=I_{\ell-1}.$
We can then take into account the values $ham({\hat{G}}_{i}),\ spi({\hat{G}}_{i}),\ |V({\hat{G}}_{i})|,$ and the sets ${\cal{P}}_{ham({\hat{G}}_{i})}$ and ${\cal{P}}_{spi({\hat{G}}_{i})}$, for each $i=1,\ldots,{\hat{n}}$. For each $B_{{\hat{{{H}}}}^{\prime}}=\{j\}$, for $j=1,\ldots,{\hat{n}}$, we can follow the approach outlined in Section 2.1 to determine the feasibility of the corresponding ILP. If feasible, according to the construction detailed in Section 2.2, it is possible to obtain a spider of ${\hat{{{G}}}}\otimes(\ell-1)$ centered in $V({\hat{G}}_{j})$.
The smallest integer $\ell$ for which the above holds determines $spi({\hat{{{G}}}})$ and provides a spider $T$ covering ${\hat{{{G}}}}\otimes(spi({\hat{{{G}}}})-1)$ with its center in $V({\hat{{{G}}}})$. The arguments outlined in Section 2.3 allow us to derive the time complexity of this computation (where, ${\hat{m}}$ represents the number of edges of ${\hat{{{H}}}}$)
$spi({\hat{{{G}}}})\;{\hat{n}}\;[O({\hat{n}}\sqrt{{\hat{n}}+{\hat{m}}})^{(1+o(1))(3{\hat{n}}+2{\hat{m}}+1)}+O({\hat{n}}({\hat{n}}+{\hat{m}}))]+O(|V({\hat{{{G}}}})|^{2}).$
It is evident that the subgraph of $T$ induced by $V({\hat{{{G}}}})$ returns the path-spider cover ${\cal{P}}_{spi({\hat{{{G}}}})}$ of ${\hat{{{G}}}}$.
Theorem 2.
Path-Spider Cover parameterized by modular-width is fixed-parameter tractable.
3.2 Computing $ham({\hat{{{G}}}})$ and ${\cal{P}}_{ham({\hat{{{G}}}})}$
Using an approach similar to the one in the proof of Fact 1, we can prove the following result.
Fact 2.
Let ${\hat{{{G}}}}$ be a graph and
$h({\hat{{{G}}}})=\min\{\ell\ |\ {\hat{{{G}}}}\otimes\ell\;\mbox{has a hamiltonian path with an end-point in $\{v_{1},\ldots,v_{\ell}\}$}\}.$
Then $ham({\hat{{{G}}}})=h({\hat{{{G}}}})$.
Proof.
We first show that $h({\hat{{{G}}}})\leq ham({\hat{{{G}}}})$. Consider $P_{1},P_{2},\ldots,P_{ham({\hat{{{G}}}})}$ the partition of ${\hat{{{H}}}}$ in disjoint paths, where $f(P_{i})$ and $s(P_{i})$ denote the first and second end-point of path $P_{i}$, respectively, for $i=1,\ldots,ham({\hat{{{G}}}})$. Consequently, the graph ${\hat{{{G}}}}\otimes ham({\hat{{{G}}}})$ contains the hamiltonian path $P$ obtained connecting vertex $v_{1}$ to $f(P_{1})$, vertex $v_{i}$ to both $s(P_{i-1})$ and $f(P_{i})$ for each for $i=2,\ldots,ham({\hat{{{G}}}})$.
Next, we establish that $ham({\hat{{{G}}}})\leq h({\hat{{{G}}}})$. Let $P$ be a hamiltonian path in ${\hat{{{G}}}}\otimes h({\hat{{{G}}}})$ with an end-point in $\{v_{1},\ldots,v_{h({\hat{{{G}}}})}\}$. Removing the vertices in $\{v_{1},\ldots,v_{h({\hat{{{G}}}})}\}$ from $P$ we obtain a partition of ${\hat{{{G}}}}$ in $h({\hat{{{G}}}})$ pairwise disjoint paths. ∎
By Fact 2, the value $ham({\hat{{{G}}}})$ is equal to the smallest positive integer $\ell$ with $1\leq\ell\leq|V({\hat{{{G}}}})|$ such that the graph
${\hat{{{G}}}}\otimes\ell$ has a hamiltonian path with an end-point in $\{v_{1},\ldots,v_{\ell}\}$. (13)
To determine if graph ${\hat{{{G}}}}\otimes\ell$ has a hamiltonian path and find it if it exists, we can follow the approach outlined in Section 2. Given that ${\hat{{{G}}}}\otimes\ell={\hat{{{H}}}}^{\prime}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}},I_{\ell})$, where $I_{\ell}$ is an independent set with $\ell$ vertices, we know that $ham(I_{\ell})=\ell$ and ${\cal{P}}_{ham(I_{\ell})}=I_{\ell}$. Therefore, with the values $ham({\hat{G}}_{i}),|V({\hat{G}}_{i})|$ and the set ${\cal{P}}_{ham({\hat{G}}_{i})}$ available for each $i=1,\ldots,{\hat{n}}$, we can construct the corresponding ILP as described in Section 2.1 setting $B_{{\hat{{{H}}}}^{\prime}}=\emptyset$ and $r={\hat{n}}+1$ (i.e., ${\hat{G}}_{r}=I_{\ell}$). If the ILP admits a solution, we can construct the hamiltonian path $P$ of ${\hat{{{G}}}}\otimes\ell={\hat{{{H}}}}^{\prime}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}},I_{\ell})$ using the method outlined in Section 2.2, choosing any vertex in ${\hat{G}}_{r}=I_{\ell}$ as end-point (i.e., root). It is important to note that all proofs and constructions detailed in Section 2.2 remains valid in this scenario (i.e., ${\hat{{{H}}}}^{\prime}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}},I_{\ell})$, $B_{{\hat{{{H}}}}^{\prime}}=\emptyset$ and ${\hat{G}}_{r}=I_{\ell}$). Furthermore, as shown in Section 2.4, it can be proved that if a hamiltonian path exists in ${\hat{{{H}}}}^{\prime}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}},I_{\ell})$ with an end-point in $I_{\ell}$, then exists a solution $(x,y)$ of the corresponding ILP (using the same arguments as in the proof of Lemma 7 with the hamiltonian path rooted at the end-point in $I_{\ell}$).
The smallest integer $\ell$ for which (13) occurs determines $ham({\hat{{{G}}}})$ and also provides the hamiltonian path $P$ of ${\hat{{{G}}}}\otimes ham({\hat{{{G}}}})$ with one end-point in $I_{ham({\hat{{{G}}}})}$. The time complexity of this computation can be derived using the arguments presented in Section 2.3. We get
$ham({\hat{{{G}}}})\;[O({\hat{n}}\sqrt{{\hat{n}}+{\hat{m}}})^{(1+o(1))(3{\hat{n}}+2{\hat{m}}+1)}+O({\hat{n}}({\hat{n}}+{\hat{m}}))]+O(|V({\hat{{{G}}}})|^{2}).$
Obviously, the subgraph of $P$ induced by $V({\hat{{{G}}}})$ will return the partition in $ham({\hat{{{G}}}})$ disjoint paths of ${\hat{{{G}}}}$, ${\cal{P}}_{ham({\hat{{{G}}}})}$.
We stress that $ham({\hat{{{G}}}})=1$ iff the graph ${\hat{{{G}}}}$ has a hamiltonian path.
Theorem 3.
Partitioning into Paths parameterized by modular-width is fixed-parameter tractable.
4 The weighted MBV  problem
Let $G=(V,E,w)$ be a weighted graph where $w(u)$ is a positive integer representing the cost of choosing $u$ as a branch vertex, for each $u\in V$. Given a spanning tree $T$ of $G$, let $B$ be the set of branch vertices in $T$. The cost of $T$ is the sum of the costs of the vertices in $B$, that is $w(T)=\sum_{u\in B}w(u)$. Denote by $w(G)$ the minimum cost of any spanning tree of $G$, the weighted version of MBV is formally defined as follows.
Minimum Cost Branch Vertices (CBV)
Instance: A weighted undirected connected graph $G=(V,E,w)$.
Goal: Find a spanning tree $T$ of $G$ having cost $w(G)$.
Trivially, in case $w(u)=1$ for each $u\in V$ then CBV  becomes the problem of determining a spanning tree having the minimum number of branch vertices (MBV).
We present an FPT algorithm for CBV parameterized by neighborhood diversity.
Recall that a connected graph $G$ of neighborhood diversity ${\tt{nd}}$, with type partition ${\cal{V}}=\{V_{1},\cdots,V_{\tt{nd}}\}$ and type graph $H$ can be seen as $G=H(G_{1},\ldots,G_{\tt{nd}})$, where $G_{i}=G[V_{i}]$ is the subgraph of $G$ induced by $V_{i}$ that is either a clique or an independent set, for each $i=1,\ldots,{\tt{nd}}$.
Lemma 8.
Let $G=H(G_{1},\ldots,G_{\tt{nd}})$ be a connected graph with type partition $,{\cal{V}}=\{V_{1},V_{2},$$\ldots,V_{\tt{nd}}\}$ and type graph $H$. For any spanning tree $T$ of $G$ with cost $w(G)$ it holds:
  • $|B\cap V_{i}|\leq 1$,
  • if $|B\cap V_{i}|=1$ then $B\cap V_{i}$ consists of a vertex having minimum cost in $V_{i}$,
for each $i=1,\ldots{\tt{nd}}$.
Proof.
Let $T$ be a spanning tree of $G$ with branch cost $w(G)$. Denote by $N_{T}(v)$ the set of neighbors of $v$ in $T$, for any vertex $v$, and by $d_{T}(v,w)$ the length of the path between $v$ and $w$ in $T$. Let $u_{i}$ be the vertex in $V_{i}$ having the minimum cost, for any $1\leq i\leq{\tt{nd}}$. We proceed by contradiction.
Assume that $|V_{i}\cap B|\geq 1$ and that no vertex $u_{i}$ of minimum cost $c(u_{i})$ is in $V_{i}\cap B$, for some $1\leq i\leq{\tt{nd}}$. Let $v$ be any vertex in $V_{i}\cap B$.
In case $G_{i}$ is an independent set, let $v^{\prime}\in N_{T}(v)$. We can modify $T$ to have a new spanning tree $T^{\prime}$, as follows: For any $x\in N_{T}(v)\setminus\{v^{\prime}\}$, substitute in $T$ the edge $\{v,x\}$ by the edge $\{u_{i},x\}$, then add the edge $\{u_{i},v^{\prime}\}$ in case it is not an edge in $T$. Hence, the set of branch vertices $B^{\prime}$ of $T^{\prime}$ is $B^{\prime}=(B\setminus\{v\})\cup\{u_{i}\}$ and $w(T^{\prime})=w(T)-w(v)+w(u_{i})<w(T)$, having a contradiction.
In case $G_{i}$ is a clique, we can modify $T$ to have a new spanning tree $T^{\prime}$, as follows: For any $x\in N_{T}(v)$, substitute in $T$ the edge $\{v,x\}$ by the edge $\{u_{i},x\}$, then add the edge $\{u_{i},v\}$ in case it is not an edge in $T$. Again, the set of branch vertices $B^{\prime}$ of $T^{\prime}$ is $B^{\prime}=(B\setminus\{v\})\cup\{u_{i}\}$ and $w(T^{\prime})=w(T)-w(v)+w(u_{i})<w(T)$, having a contradiction.
Assume now that $u_{i}\in V_{i}\cap B$ and $|V_{i}\cap B|\geq 2$. Let $v\in V_{i}\cap B$ with $v\neq u_{i}$.
Consider first the case in which $G_{i}$ is an independent set. Consider the path connecting $u_{i}$ and $v$ in $T$, say $u_{i},\ldots,v^{\prime\prime},v^{\prime},v$ We can modify $T$ to have a new spanning tree $T^{\prime}$, as follows: For any $x\in N_{T}(v)\setminus\{v^{\prime}\}$, substitute in $T$ the edge $\{v,x\}$ by the edge $\{u_{i},x\}$, then, in case $d_{T}(u_{i},v)\geq 3$, substitute in $T$ the edge $\{v^{\prime\prime},v^{\prime}\}$ by the edge $\{u_{i},v^{\prime}\}$. Hence, the set of branch vertices $B^{\prime}$ of $T^{\prime}$ is $B^{\prime}=B\setminus\{v\}$ and $w(T^{\prime})<w(T)$, having a contradiction.
Consider the case in which $G_{i}$ is a clique. We can modify $T$ to have a new spanning tree $T^{\prime}$, as follows: For any $x\in N_{T}(v)\setminus\{u_{i}\}$, substitute in $T$ the edge $\{v,x\}$ by the edge $\{u_{i},x\}$, then, in case $d_{T}(u_{i},v)\geq 2$, add the edge $\{u_{i},v\}$. Again, the set of branch vertices $B^{\prime}$ of $T^{\prime}$ is $B^{\prime}=B\setminus\{v\}$ and $w(T^{\prime})<w(T)$, having a contradiction. ∎
By using Lemma 8, we design an algorithm that constructs a spanning tree of $G$ with branch cost $w(G)$, having at most one branch vertex in each type set $V_{i}$ of the type partition ${\cal{V}}$ of $G$, and where, if $V_{i}$ contains a branch vertex then it is the vertex $u_{i}$ with minimum cost in $V_{i}$.
Our algorithm uses a simplified version of the algorithm presented in Section 2. We know that for each $i=1,\ldots,{\tt{nd}}$ the subgraph $G_{i}=G[V_{i}]$ induces either a clique or an independent set. As a consequence, we have that:
  • If $G_{i}$ induces a clique then
    $spi(G_{i})=1$ and the path-spider cover of $G_{i}$, ${\cal{P}}_{spi(G_{i})}$, is simply a star graph. We assume that the center of such a star is the vertex $u_{i}$ with minimum cost in $V_{i}$.
    $ham(G_{i})=1$ and the partition of $G_{i}$, ${\cal{P}}_{ham(G_{i})}$, consists of one path going through all the vertices in $V_{i}$.
  • If $G_{i}$ induces an independent set then
    $spi(G_{i})=ham(G_{i})=|V_{i}|$ and ${\cal{P}}_{spi(G_{i})}={\cal{P}}_{ham(G_{i})}=V_{i}$.
Hence, consider all the subsets of $\{1,\cdots,{\tt{nd}}\}$ (with $|B_{H}|\geq 1$) arranged in ascending order of cost. For each non empty $B_{H}\subseteq\{1,\cdots,{\tt{nd}}\}$, if we choose two vertices $r\in V(H)=\{1,\ldots,{\tt{nd}}\}$ and $s\notin\{1,\ldots,{\tt{nd}}\}$, and consider directed graph ${A_{B_{{H}}}}=\{(s,r)\}\cup\{(i,j),(j,i)\ |\ \{i,j\}\in E(H)\}$, the ILP in Section 2.1 becomes:
$x_{sr}=1$ (14)
$\sum_{j:(j,i)\in{A_{B_{{H}}}}}x_{ji}\leq|V_{i}|\qquad\qquad\qquad\forall i\in\{1,\ldots,{\tt{nd}}\}\mbox{\ s.t. $G_{i}$ is a clique}$ (15)
[…]
$\sum_{\ell:(i,\ell)\in{A_{B_{{H}}}}}x_{i\ell}-\sum_{j:(j,i)\in{A_{B_{{H}}}}}x_{ji}\leq 0\qquad\ \forall i\in\{1,\ldots,{\tt{nd}}\}\setminus B_{H}$ (17)
$y_{sr}={\tt{nd}}$ (18)
$\sum_{j:(j,i)\in{A_{B_{{H}}}}}y_{ji}-\sum_{\ell:(i,\ell)\in{A_{B_{{H}}}}}y_{i\ell}=1\ \qquad\forall i\in\{1,\ldots,{\tt{nd}}\}$ (19)
$y_{ij}\leq{\tt{nd}}\ x_{ij}\qquad\qquad\qquad\ \ \ \ \forall(i,j)\in{A_{B_{{H}}}}$ (20)
$y_{ij},x_{ij}\in\mathbb{N}\qquad\qquad\qquad\ \ \ \ \ \ \forall(i,j)\in{A_{B_{{H}}}}$ (21)
Whenever a feasible solution of the above ILP exists, the construction of the spanning tree $T$ of $G$ with branch vertices $B=\{u_{i}\ |\ i\in B_{H}\}$ follows the lines given in Section 2.2 and algorithm TREE, once ${\cal{P}}_{spi(G_{i})}$ and ${\cal{P}}_{ham(G_{i})}$, for $i\in[{\tt{nd}}]$, are as described above.
The time complexity of the algorithm can be easily derived by considering that: The ILP is solved at most for each $B_{H}\subseteq\{1,\cdots,{\tt{nd}}\}$ and, that it has at most $q={\tt{nd}}^{2}+3{\tt{nd}}+2$ constraints, at most $p=2({\tt{nd}}^{2}+1)$ variables and $\Delta={\tt{nd}}$; the algorithm TREE takes time $O(V(G)^{2})$. By the results of Jansen and Rohwedderb [29], overall the algorithm requires time
$2^{\tt{nd}}(O({\tt{nd}}^{2})^{(1+o(1))({\tt{nd}}^{2}+3{\tt{nd}}+2)}+O({\tt{nd}}^{4}))+O(n^{2}).$
It is easy to conclude that the spanning tree of $G$ with the minimum branch cost $w(G)$ is the tree that can be obtained by using the set ${B_{{H}}}\subseteq\{1,\ldots,{\tt{nd}}\}$ having the least branch cost $w(B_{H})$ for which the ILP admits a solution. Hence, we have the following result.
Theorem 4.
Minimum Cost Branch Vertices parameterized by neighborhood diversity is fixed-parameter tractable.
5 Conclusions
We investigated the parameterized complexity of the Minimum Branch Vertices problem when parameterized by the modular width of the input graph, presenting an FPT algorithm. Additionally, we designed FPT algorithms for both the Path-Spider Cover and Partitioning into Paths problems, parameterized by modular width. Furthermore, we presented an FPT algorithm parameterized by neighborhood diversity for the weighted version of the MBV problem, the Minimum Cost Branch Vertices problem. The parameterized complexity of the Minimum Cost Branch Vertices problem remains an open question when parameterized by modular width.

### v3

@@
-
-Spanning Trees Minimizing Branching Costs ***An extended abstract of some of the results of this paper appeared in [23].
-[…]
-Department of Computer Science
-University of Salerno
-Italy
-{lgargano, arescigno}@unisa.it
+Spanning Trees Minimizing Branching Costs
+Abstract
 The Minimum Branch Vertices Spanning Tree problem aims to find a spanning tree $T$ in a given graph $G$ with the fewest branch vertices, defined as vertices with a degree three or more in $T$.
@@
 Additionally, in cases where each vertex in the input graph has an associated cost for serving as a branch vertex, we prove that the problem of finding a spanning tree with the minimum branch cost (i.e., minimizing the sum of the costs of branch vertices) is FPT with respect to neighborhood diversity.
-Keywords: Spanning tree, Fixed parameterized algorithms, Modular-width, Neighborhood diversity
-Declarations
-  • Funding: This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU.
-  • Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.
+keywords:
+Spanning tree, Fixed parameterized algorithms, Modular-width, Neighborhood diversity
 1 Introduction
@@
 The problem of determining a spanning tree with a limited number of branch vertices, though fundamentally a theoretical question, originated from addressing challenges in wavelength-division multiplexing (WDM) technology within optical networks.
-WDM is a highly effective method for utilizing the full bandwidth potential of optical fibers, thereby meeting the high bandwidth demands of the Internet [28].
+WDM is a highly effective method for utilizing the full bandwidth potential of optical fibers, thereby meeting the high bandwidth demands of the Internet He et al.
+(2002).
 Multicasting involves the simultaneous transmission of information from a single source to multiple destinations.
@@
 Due to the high cost of light-splitting switches, it is essential to minimize the number of nodes equipped with these devices.
-This need gives rise to the Minimum Branch Vertices (MBV) problem, which seeks to reduce the number of light-splitting switches in a light tree[20].
+This need gives rise to the Minimum Branch Vertices (MBV) problem, which seeks to reduce the number of light-splitting switches in a light tree Gargano et al.
+(2004).
 In Cognitive Radio Networks, as well as in 5G technologies that operate across a broad spectrum of frequencies, managing the switching costs associated with transitions between different service providers is of utmost importance.
-This is essential not only for minimizing delays but also for optimizing energy consumption [27, 37].
+This is essential not only for minimizing delays but also for optimizing energy consumption Gozupek et al.
+(2013); Shami and Rasti (2016).
 The Minimum Branch Vertices (MBV) problem has been extensively studied from both algorithmic and graph-theoretic perspectives.
@@
 Gargano et al.
-[21] proved that determining whether a graph $G$ has a spanning tree with at most $k$ branch vertices is NP-complete, for each $k\geq 1$, even in cubic graphs.
+(2002) proved that determining whether a graph $G$ has a spanning tree with at most $k$ branch vertices is NP-complete, for each $k\geq 1$, even in cubic graphs.
 In the same paper, they provided an algorithm that finds a spanning tree with one branch vertex if every set of three independent vertices in $G$ has a degree sum of at least $|V(G)|-1$.
-Salamon [36] developed an algorithm that finds a spanning tree with $O(\log|V(G)|)$ branch vertices for graphs where each vertex has a degree of $\Omega(n)$; moreover, an approximation factor better than $O(\log|V(G)|)$ would imply that $NP\subseteq DTIME(n^{O(\log\log n)})$.
-Sufficient conditions for a connected claw-free graph to have a spanning tree with $k$ branch vertices are presented in [34].
-Integer linear formulations of the MBV problem and its variants are discussed in [9, 10, 11], including different relaxations and numerical comparisons of these relaxations.
-Hybrid integer linear programs for MBV are considered and solved with branch-and-cut algorithms in [39].
-Decomposition methods for solving the MBV problem are explored in [31, 33, 35], while other heuristics are presented in [32, 38, 40].
-A complementary formulation called the Maximum Path-Node Spanning Tree (MPN), which aims to maximize the number of vertices with a degree of at most two, is investigated in [12].
+Salamon (2005) developed an algorithm that finds a spanning tree with $O(\log|V(G)|)$ branch vertices for graphs where each vertex has a degree of $\Omega(n)$; moreover, an approximation factor better than $O(\log|V(G)|)$ would imply that $NP\subseteq DTIME(n^{O(\log\log n)})$.
+Sufficient conditions for a connected claw-free graph to have a spanning tree with $k$ branch vertices are presented in Matsuda et al.
+(2014).
+Integer linear formulations of the MBV problem and its variants are discussed in Carrabs et al.
+(2013); Cerrone et al.
+(2014); Cerulli et al.
+(2009), including different relaxations and numerical comparisons of these relaxations.
+Hybrid integer linear programs for MBV are considered and solved with branch-and-cut algorithms in Silvestri et al.
+(2017).
+Decomposition methods for solving the MBV problem are explored in Landete et al.
+(2017); Melo et al.
+(2016); Rossi et al.
+(2014), while other heuristics are presented in Marin (2015); Silva et al.
+(2014); Sundar et al.
+(2012).
+A complementary formulation called the Maximum Path-Node Spanning Tree (MPN), which aims to maximize the number of vertices with a degree of at most two, is investigated in Chimani and Spoerhase (2015).
 The authors prove that MPN is APX-hard and provide an approximation algorithm with a ratio of 6/11.
-Related gathering processes are considered in [5, 6, 7, 24, 25].
+Related gathering processes are considered in Bermond et al.
+(2008, 2010, 2013); Gargano and Rescigno (2015, 2009).
 1.1 Graph Partitioning.
@@
 We will denote by ${\cal{P}}_{ham(G)}$ a partition of $G$ into $ham(G)$ paths.
-Notice that the Partitioning into Paths problem was already considered in [18]; however as originally defined, it only asks for the value $ham(G)$, while we ask for the actual path partitioning of $G$.
-Recently, the PP problem and some its variants have been considered in [16].
+Notice that the Partitioning into Paths problem was already considered in Gajarský et al.
+(2013); however as originally defined, it only asks for the value $ham(G)$, while we ask for the actual path partitioning of $G$.
+Recently, the PP problem and some its variants have been considered in Fernau et al.
+(2023).
 When referring to a path $P$ in $G$, we denote its end-points as $f(P)$ and $s(P)$, distinguishing them as the first and the second end-point of $P$, respectively.
@@
 For the weighted MBV problem we present an FPT algorithm parameterized by neighborhood diversity.
-An FPT algorithm for the case of uniform cost was given in [22].
+An FPT algorithm for the case of uniform cost was given in Gargano and Rescigno (2023b).
 1.3 Parameterized Complexity.
@@
 A problem with input size $n$ and parameter $p$ is called fixed parameter tractable (FPT) if it can be solved in time $f(p)\cdot n^{c}$, where $f$ is a computable function only depending on $p$ and $c$ is a constant.
-It was recently proven that MBV is FPT when parameterized by treewidth [4].
-The algorithm presented in [4] runs in time $O(4^{2{\tt tw}}{\tt tw}^{4{\tt tw}+1}({\tt tw}+1)\;n)$, where tw is the treewidth of the input graph.
-On the other hand, it was shown in [17] that the problem is $W[1]$-hard when parameterized by clique-width.
-Specifically, in [17] it was proven that the (MBV special case) hamiltonian path problem is $W[1]$-hard when parameterized by clique-width.
+It was recently proven that MBV is FPT when parameterized by treewidth Baste and Watel (2022).
+The algorithm presented in Baste and Watel (2022) runs in time $O(4^{2{\tt tw}}{\tt tw}^{4{\tt tw}+1}({\tt tw}+1)\;n)$, where tw is the treewidth of the input graph.
+On the other hand, it was shown in Fomin et al.
+(2009) that the problem is $W[1]$-hard when parameterized by clique-width.
+Specifically, in Fomin et al.
+(2009) it was proven that the (MBV special case) hamiltonian path problem is $W[1]$-hard when parameterized by clique-width.
 In this paper, we are interested in assessing the complexity of MBV when parameterized by modular width and, in the weighted case, by neighborhood diversity.
@@
 Dashed arrows denote that the generalization may exponentially increase the parameter.
-Modular-width. Modular-width was introduced in [18] and is defined as follows.
+Modular-width. Modular-width was introduced in Gajarský et al. (2013) and is defined as follows.
 Consider graphs that can be obtained from an algebraic expression that uses the following operations:
@@
 The modular-width of a graph ${{G}}$, denoted ${\tt{mw}}({{G}})$, is the least integer $m$ such that ${{G}}$ can be obtained from such an algebraic expression of width at most $m$.
-We recall that an algebraic expression of width ${\tt{mw}}({{H}})$ can be constructed in linear time [41].
+We recall that an algebraic expression of width ${\tt{mw}}({{H}})$ can be constructed in linear time Tedder et al. (2008).
 Given a graph $G=(V,E)$, denote by $N(u)$ the set of neighbors of vertex $u\in V$.
@@
 Neighborhood diversity.
-The neighborhood diversity parameter, was first introduced by Lampis in [30].
-It has then received much attention [1, 2, 8, 13, 14, 15, 19, 24, 26, 42], also due to the fact that, contrary to other popular parameters, it is computable in linear time.
+The neighborhood diversity parameter, was first introduced by Lampis in Lampis (2012).
+It has then received much attention Agrawal et al.
+(2020); Araujo et al.
+(2018); Bhyravarapu and Reddy (2022); Cordasco et al.
+(2018, 2020, 2022); Ganian (2012); Gargano and Rescigno (2015); Gavenciak et al.
+(2022); Toufar et al.
+(2019), also due to the fact that, contrary to other popular parameters, it is computable in linear time.
 Given a graph $G=(V,E)$, two vertices $u,v\in V$ are said to have the same type if $N(v)\setminus\{u\}=N(u)\setminus\{v\}$.
@@
 By repeating steps a) and b), one can construct the desired spanning tree of ${{G}}$ with at most one branch vertex in each $V(G_{i})$.
-∎
 In the remaining part of this section, we present an algorithm that computes a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $b({{G}})$ branch vertices, if $b({{G}})>0$.
 In Section 3.2 we deal with the case $b({{G}})=0$, that is, we show how to find a Hamiltonian path of ${{G}}$, if one exists.
-The work in [18] focuses on verifying the existence of a Hamiltonian path in a graph; the proofs therein implicitly give a method for constructing such a path.
+The work in Gajarský et al.
+(2013) focuses on verifying the existence of a Hamiltonian path in a graph; the proofs therein implicitly give a method for constructing such a path.
 We recall such a construction and show its relationship as a special case of our algorithm.
@@
 The process continues iteratively as long as there are unexplored vertices, i.e., $V({{G}})\setminus Ex\neq\emptyset$.
-The EXPLORE$(u)$ procedure initiates exploration from $u=f(P^{j}_{k})$ along with the entire path or spider $P^{j}_{k}$ Assume that when in the algorithm $P\in{\cal{P}}^{j}$ is explored (i.e., $Ex=Ex\cup P$), the parent function $\pi$ is updated.
+The EXPLORE$(u)$ procedure initiates exploration from $u=f(P^{j}_{k})$ along with the entire path or spider $P^{j}_{k}$  [^1] Assume that when in the algorithm $P\in{\cal{P}}^{j}$ is explored (i.e., $Ex=Ex\cup P$), the parent function $\pi$ is updated.
 For path, $\pi$ is set from $f(P)$ to $s(P)$, and for spiders, it is set from the center $f(P)=s(P)$ to the leaves..
@@
   • if a solution exists for the current set ${B_{{H}}}$, use algorithm TREE to construct a spanning tree of ${{G}}={{H}}(G_{1},\ldots,G_{{n}})$ with $|{B_{{H}}}|$ branch vertices.
-Jansen and Rohwedderb [29] have recently showed that the time needed to find a feasible solution of an ILP with $p$ integer variables and $q$ constraints is $O(\sqrt{q}\Delta)^{(1+o(1))q}+O(qp)$, where $\Delta$ is the largest absolute value of any coefficient in the ILP.
+Jansen and Rohwedderb (2023) have recently showed that the time needed to find a feasible solution of an ILP with $p$ integer variables and $q$ constraints is $O(\sqrt{q}\Delta)^{(1+o(1))q}+O(qp)$, where $\Delta$ is the largest absolute value of any coefficient in the ILP.
 Denoted by ${{m}}$ the number of edges of ${{H}}$, our ILP has $q=3{{n}}+2{{m}}+1$ constraints, $p=2({{m}}+1)$ variables and $\Delta={{n}}$.
@@
 3 The triple and partition computation
-Following [18], we use a bottom-up dynamic programming approach along the parse-tree of an algebraic expression describing ${{G}}$, to compute for every internal vertex ${\hat{{{G}}}}={\hat{{{H}}}}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}})$ of the parse tree a record of data, using those already computed for its children.
+Following Gajarský et al.
+(2013), we use a bottom-up dynamic programming approach along the parse-tree of an algebraic expression describing ${{G}}$, to compute for every internal vertex ${\hat{{{G}}}}={\hat{{{H}}}}({\hat{G}}_{1},\ldots,{\hat{G}}_{\hat{n}})$ of the parse tree a record of data, using those already computed for its children.
 Specifically, given the triple $(ham({\hat{G}}_{i}),spi({\hat{G}}_{i}),|V({\hat{G}}_{i})|)$,  ${\cal{P}}_{ham({\hat{G}}_{i})}$ and ${\cal{P}}_{spi({\hat{G}}_{i})}$, for each $i=1,\ldots,{\hat{n}}$, we need to compute
@@
 Goal: Find a spanning tree $T$ of $G$ having cost $w(G)$.
-Trivially, in case $w(u)=1$ for each $u\in V$ then CBV  becomes the problem of determining a spanning tree having the minimum number of branch vertices (MBV).
+. Trivially, in case $w(u)=1$ for each $u\in V$ then CBV  becomes the problem of determining a spanning tree having the minimum number of branch vertices (MBV).
 We present an FPT algorithm for CBV parameterized by neighborhood diversity.
@@
 The time complexity of the algorithm can be easily derived by considering that: The ILP is solved at most for each $B_{H}\subseteq\{1,\cdots,{\tt{nd}}\}$ and, that it has at most $q={\tt{nd}}^{2}+3{\tt{nd}}+2$ constraints, at most $p=2({\tt{nd}}^{2}+1)$ variables and $\Delta={\tt{nd}}$; the algorithm TREE takes time $O(V(G)^{2})$.
-By the results of Jansen and Rohwedderb [29], overall the algorithm requires time
+By the results of Jansen and Rohwedderb Jansen and Rohwedderb (2023), overall the algorithm requires time
 $2^{\tt{nd}}(O({\tt{nd}}^{2})^{(1+o(1))({\tt{nd}}^{2}+3{\tt{nd}}+2)}+O({\tt{nd}}^{4}))+O(n^{2}).$